### v3

# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

Abstract
Prior-data fitted networks (PFNs) have emerged as promising foundation models for prediction from tabular datasets, achieving state-of-the-art performance on small to moderate data sizes without tuning. While PFNs are motivated by Bayesian ideas, they do not provide any uncertainty quantification for predictive means, quantiles, or similar quantities. We propose a principled and efficient sampling procedure to construct Bayesian posteriors for such estimates based on martingale posteriors, and prove its convergence. Several simulated and real-world data examples showcase the uncertainty quantification of our method in inference applications.
Machine Learning, ICML
1 Introduction
Prior-data fitted networks (PFNs) are foundation models (Hollmann et al., 2023; Müller et al., 2022) that allow for in-context learning, i.e., the ability to learn at inference time without any parameter updates (Garg et al., 2022). TabPFN, a transformer pre-trained on synthetic data for in-context learning on tabular datasets, has recently attracted a lot of interest. TabPFN (Hollmann et al., 2023, 2025) and extensions such as TuneTables (Feuer et al., 2024) and LocalPFN (Thomas et al., 2024) have been shown to achieve state-of-the-art performance on tabular benchmarks by pre-training on purely synthetic data. Since PFNs and extensions learn in-context, there is no need for further model (fine-)tuning on the inference task.
Recent extensions of PFNs allow their applicability to large datasets (Feuer et al., 2024), the use of PFN “priors” for latent variable models (Reuter et al., 2025), and simultaneously minimizing bias and variance to improve their performance (Liu and Ye, 2025). PFNs are also related to simulation-based inference and amortized inference but have slightly different goals and do not amortize across a single but multiple datasets (Reuter et al., 2025). While introduced as a Bayesian method and approximation to the posterior predictive, PFNs can also be interpreted as pre-tuned untrained predictors (Nagler, 2023). This also relates to the question of what uncertainty PFN models can provide.
PFNs approximate the posterior predictive distribution for the label given some feature values. Despite the name, this only yields point estimates of the most relevant predictive quantities, such as the conditional mean or quantiles. Due to the complex nature of PFNs, it is difficult to assess the uncertainty of these point estimates.
We propose a principled and efficient method to construct Bayesian posteriors for such estimates using the idea of Martingale Posteriors (MPs) proposed in Fong et al. (2023). In particular:
  1. 1.
    We introduce a new extension of the MP framework for inference of predictive quantities conditional on a specific feature value $x$.
  2. 2.
    We propose an efficient, nonparametric resampling scheme yielding an approximate posterior for the point estimates derived from a PFN, and prove its convergence.
  3. 3.
    We extend existing learning rate schedules and discuss the role of contraction rates in light of inconsistent PFN estimators.
  4. 4.
    We illustrate the new method in several simulated and real-world data applications.
  5. 5.
    We perform a variety of ablation studies providing insights into the driving factors of our proposal’s efficacy and failure modes of alternative methods.
Our work provides an essential tool for principled inference with the exponentially popular PFN methods and overcomes limitations of previous methods that struggle with complex datasets or PFN’s missing martingale property (cf. Figure 1). A brief discussion of potential extensions to large language models is provided in Section 5.
Figure 1: Comparison of estimated median function (orange line) and credible intervals (shaded blue area) for a non-linear feature effect (dashed blue line) in a Bayesian additive model based on different methods (facets). Numbers in brackets after each method state the marginal miscoverage, i.e., the absolute difference between the method’s average coverage and the nominal level of 90% (smaller is better). Conditional coverages for individual $x_{1}$ values are provided in the inset plots with the nominal level marked as a gray dashed line. “Optimal” corresponds to the posterior induced by the chosen prior and likelihood, which in this case can be analytically derived (see Section 4.3 for details). Its miscoverage is 0.018 (which is plausible, being in a small data setting) and very close to our method.
2 BACKGROUND
We consider a tabular prediction task with labels $y\in\mathbb{R}$ and features $x\in\mathbb{R}^{d}$ drawn from a joint distribution $P$. A typical problem in such tasks is to estimate predictive quantities such as conditional means $\mathbb{E}[y|x]$, conditional probabilities $P(y|x)$, or conditional quantiles $P^{-1}(\alpha|x)$. Because the true distribution $P$ is unknown and only a finite amount of data $\mathcal{D}_{n}=(y_{i},x_{i})_{i=1}^{n}$ is available, estimates of such quantities bear some uncertainty. Our goal is to quantify this uncertainty.
2.1 Prior-Data Fitted Networks
Prior-data fitted networks are foundation models trained to approximate the posterior predictive density $\operatorname{PPD}(y|x)=p(y|x,\mathcal{D}_{n})$, which quantifies the likelihood of observing label $y$ given that the feature is $x$ and $\mathcal{D}_{n}$ has been observed. The PPD is a Bayesian concept and implicitly involves a prior over the distributions $P$ that could have generated the data. To approximate the PPD with a PFN, a deep neural network—typically a transformer—is pre-trained on simulated datasets with diverse characteristics. After pre-training, the network weights are fixed, and the approximate PPD for a new training set can be computed through a single forward pass without additional training or tuning.
The PPD quantifies uncertainty about the label $y$. However, it mixes the aleatoric and epistemic components of uncertainty (Hüllermeier and Waegeman, 2021). From the PPD alone, it is impossible to disentangle these parts. Consequently, PFNs do not provide uncertainty estimates for predictive summaries, such as conditional mean, probabilities, and quantiles.
2.2 Bayesian Inference
In classical Bayesian inference, the set of possible distributions $P=P_{\theta}$ is indexed by some parameter $\theta$. A prior distribution $\pi(\theta)$ is elicited to quantify our beliefs about the likelihood of the possible values of $\theta$ before seeing any data. After observing $\mathcal{D}_{n}$, this belief is updated to a posterior $\pi(\theta|\mathcal{D}_{n})$ of the parameter $\theta$ given the data. For predictive inference, the PPD can be computed as
$\operatorname{PPD}(y|x)=\int p_{\theta}(y|x)\pi(\theta|\mathcal{D}_{n})\,d\theta.$
In contrast to the PPD, the posterior $\pi(\theta|\mathcal{D}_{n})$ also quantifies uncertainty for other interest quantities. For example, the posterior distribution for the conditional mean $\mu(x)=\int y\,p_{\theta}(y|x)dy$ is given by
$\Pi(\mu(x)\in A)=\int\bm{1}\left\{\int y\,p_{\theta}(y|x)dy\in A\right\}\pi(\theta|\mathcal{D}_{n})\,\,d\theta,$
for any $A\subseteq\mathbb{R}$. PFNs neither provide an explicit model for $p_{\theta}$ nor an explicit prior $\pi(\theta)$, although both may be implicit in the PPD. The following shows how Bayesian posterior inference can be approached when only the PPD is available.
2.3 Martingale Posteriors
Martingale posteriors were recently introduced by Fong et al. (2023) as a new method for Bayesian uncertainty quantification. Its core idea is to reverse the direction of the Bayesian inference. In classical Bayesian inference, the posterior is derived from a prior and likelihood, which then implicitly leads to the PPD. MP inference starts from the PPD and leaves the prior $\pi(\theta)$ implicit. An appropriate sampling scheme and Doob’s theorem then allow us to derive posteriors for virtually all quantities of interest (e.g., the conditional mean $\mu(x)$).
To simplify our outline of the approach, consider the case where there are no features, and we are interested in unconditional inference. An extension to our predictive inference setting will be made explicit in Section 3.1. Suppose we have observed data $y_{1:n}=(y_{1},\dots,y_{n})$.
The MP approach involves iteratively sampling
$y_{n+1}\sim p(\cdot|y_{1:n}),\quad y_{n+2}\sim p(\cdot|y_{1:(n+1)}),$ (1)
$y_{n+3}\sim p(\cdot|y_{1:(n+2)}),\quad\dots$ (2)
NN times, which yields a sample $y_{(n+1):(n+N)}$ drawn from the predictive joint distribution
$p(y_{(n+1):(n+N)}|y_{1:n})=\prod_{i=1}^{N}p(y_{n+i}|y_{1:(n+i-1)}).$
Observe, however, that the samples are neither independent nor identically distributed. As a consequence, the long-run empirical distribution of the obtained sample,
$F_{\infty}(y)=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\bm{1}(y_{n+i}\leq y),$
is a random function and comes out differently whenever the sampling procedure is repeated. Denote by $\Pi(F_{\infty}|\mathcal{D}_{n})$ the distribution of this function (which depends on the data $\mathcal{D}_{n}$ we start with). For any parameter $\theta=\theta(P)$ of interest, the martingale posterior is now given as
$\Pi(\theta\in A|\mathcal{D}_{n})=\int\bm{1}\{\theta(F_{\infty})\}d\Pi(F_{\infty}|\mathcal{D}_{n}),$
where $A$ is any Borel set on the space where the parameter $\theta$ lives. Furthermore, Doob’s theorem (Doob, 1949) implies that $\Pi(\theta|\mathcal{D}_{n})$ coincides with the classical Bayes posterior for the prior $\pi(\theta)$ implicit in the PPD (Fong et al., 2023).
3 Efficient Martingale Posteriors for Prior-Data Fitted Networks
Martingale posteriors allow for Bayesian inference directly from the PPD. PFNs approximate the PPD, so using PFNs to construct a martingale posterior seems natural. However, there are two problems. First, modern PFNs are based on transformer architectures that require $\Omega(n^{2})$ operations for a forward pass on a training set size of $n$. Iteratively computing $p(y|y_{1:(n+k)})$ for $k=1,\dots,N$ thus has complexity $\Omega(N^{3})$, which is prohibitive. Second, Falck et al. (2024) found that modern transformer-based LLMs substantially deviate from the martingale property
$\mathbb{E}[p(y|y_{1:(n+k)})|y_{1:n}]=p(y|y_{1:n}).$
Also TabPFN fails in this regard.
Figure 2: PFN’s initial CDF (x-axis) vs. its average CDF of subsequent iterations (y-axis).
LABEL:{fig:martprop} (experiment description in Section 4.1) shows a QQ-plot of the average CDF resulting from PFN’s PPD at iteration 0 vs. the CDF after $k$ forward iterations of the sampling procedure. The martingale property is satisfied when this distribution remains constant, i.e., all lines lie on the $x=y$ diagonal. However, we see an apparent change in the distribution over the different sampling horizons. This reinforces that PFNs approximate Bayesian inference only in a weak sense, even though they were conceived from Bayesian ideas (Hollmann et al., 2023; Müller et al., 2021).
Without the martingale property, the MP sampling procedure leads to meaningless results. Instead, we propose to use the PPD implied by the PFN only as a starting point for the sampling scheme. This PPD is then iteratively updated using the nonparametric Gaussian copula approach of Fong et al. (2023), which ensures the martingale property. The resulting martingale posterior effectively treats the PFN output as an informed prior, without requiring the PFN to provide coherent posterior updates.
3.1 Martingale Posteriors for Conditional Inference
We extend the unconditional sampling scheme outlined in the previous section to the conditional inference setting. Fong et al. (2023) already proposed one such extension. Their scheme involves forward sampling of the features $x_{(n+1):(n+N)}$. The distribution of the features isn’t of primary interest but significantly complicates the sampling procedure, which the authors resolved through heuristic simplifications. In contrast to Fong et al. (2023), we propose to sample only the labels $y_{(n+1):(n+N)}$ conditional on the event that $x_{n+k}=x$, for a fixed value of $x$ and all $k=1,\dots,N$.
Specifically, our goal is to simulate data from the distribution
$p(y_{(n+1):(n+N)}|x_{(n+1):(n+N)}=x,\mathcal{D}_{n}).$
Set $x_{n+k}=x$ for all $k\geq 1$, and define
$p_{k}(y)=p(y_{n+k+1}|y_{1:(n+k)},x_{1:(n+k)}),$
and $P_{k}$ as the corresponding CDF. Applying Bayes’ rule recursively gives
$p(y_{(n+1):(n+N)}|x_{(n+1):(n+N)}=x,\mathcal{D}_{n})=\prod_{k=n}^{N-1}p_{k}(y_{k+1}),$
which suggests that we can iteratively sample
$y_{n+1}\sim P_{0},\quad y_{n+2}\sim P_{1},\quad y_{n+3}\sim P_{2},\quad\dots.$
Denote the long-run empirical distribution of the obtained sample by $F_{\infty,x}=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\bm{1}(y_{n+i}\leq y)$, which is again a random function, even in the limit. Repeating the iterative sampling procedure gives us its distribution $\Pi(F_{\infty,x}|\mathcal{D}_{n})$. For any conditional parameter $\theta(x)=\theta(P(\cdot\,|x))$ of interest, the martingale posterior is now given as
$\Pi(\theta(x)\in A|\mathcal{D}_{n})=\int\bm{1}\{\theta(F_{\infty,x})\}d\Pi(F_{\infty,x}|\mathcal{D}_{n}).$
Common examples of the parameter $\theta(x)$ are the conditional mean $\theta(x)=\int y\,dP(y|x)dy$ or a conditional $\varphi$-quantile $\theta(x)=P^{-1}(\varphi|x)$.
3.2 Efficient PPD Updates Based on the Gaussian Copula
Observe that $p_{0}(y)=p(y|x,\mathcal{D}_{n})$ is the PPD approximated by the PFN $g_{\theta}$. However, the following update distributions $p_{1},p_{2},\dots$ are generally intractable. To alleviate this, Fong et al. (2023) proposed a computationally efficient, nonparametric method based on Dirichlet Process Mixture Models (DPMMs) and a copula decomposition of the conditional $p_{k}$. Specifically, we set
$P_{k}(y)=$ $(1-\alpha_{n+k})P_{k-1}(y)$ (3)
$+\alpha_{n+k}H_{\rho}(P_{k-1}(y),P_{k-1}(y_{n+k})),$
where $P_{k}$ is the CDF corresponding to $p_{k}$, $\alpha_{i}=\left(2-{i}^{-1}\right)(i+1)^{-1}$,
$H_{\rho}(u,v)=\Phi\left(\frac{\Phi^{-1}(u)-\rho\Phi^{-1}(v)}{\sqrt{1-\rho^{2}}}\right),$
and $\Phi$ is the standard normal cumulative distribution function. The method has a hyperparameter $\rho$, corresponding to a bandwidth that smoothes the updates. Fong et al. (2023) proposed to tune this by maximizing the likelihood of the updated densities over the observed data. To increase the alignment of the updates with the PFN baseline, we simulate a new sample from the PFN and optimize the bandwidth on the simulated data. More details can be found in Section B.1
3.3 Flexible Learning Rate Schedules
Fong et al. (2023) suggested the learning rate $\alpha_{i}=(2-i^{-1})(i+1)^{-1}$ for the unconditional context, where distribution functions can be estimated with convergence rate $n^{-1/2}$. As we shall see in Proposition 3.3 below, the learning rate $\alpha_{i}$ determines the contraction rate of the posterior. Our setting, however, is conditional on $x\in\mathbb{R}^{d}$, where nonparametric methods such as TabPFN must converge slower than $n^{-1/2}$, with rates depending on the dimension and smoothness of the estimation problem (see, e.g., Stone, 1982).
To accommodate such settings, we allow for more flexible learning rate schedules that can be calibrated to match the behavior of the PFN. Specifically, we parametrize the learning rate as
$\alpha_{n+k}=C(n+k+1)^{-\beta},\quad C\in(0,\infty),\beta\in(0,1].$
where $C$ and $\beta$ should be chosen such that the behavior of the update (3) roughly matches the behavior of the PFN. In our experiments, we implement the following schedules:
  • default: $C=2$, $\beta=1$ as in (Fong et al., 2023),
  • type 1: $C=10^{-(d-1)/(d+4)}$, $\beta=1/2+d/(d+4)$,
  • type 2: $C=10^{-2d/(d+4)}$, $\beta=d/(d+4)$.
The constants $C$ are set to normalize the schedules to a comparable starting scales. We refer to our discussion after Proposition 3.3 in the next section for a motivation of the exponent $\beta$, and to Section 4.4 for empirical evaluation.
3.4 Theoretical Properties
Despite the simplicity of the updates given in (3), they provide essential theoretical guarantees. The following is a direct consequence of Theorem 3 by Fong et al. (2023).
Proposition 3.1.
It holds $(y_{N+1},y_{N+2},\dots)\to_{d}(z_{1},z_{2},\dots)$ as $N\to\infty$ where $(z_{1},z_{2},\dots)$ has an exchangeable distribution.
By de Finetti’s theorem (e.g., Theorem 1.49 in Schervish, 2012), it then follows that there is a random variable $\Theta$ such that $(z_{1},z_{2},\dots)$ is conditionally iid given $\Theta$. The distribution of $\Theta$ can be interpreted as an implicit prior. In our setting, this prior depends both on the initial PPD implied by the PFN and the Gaussian copula updates specified by (3). In particular, the following result follows from Proposition 3.1 above and Theorem 2.2 of Berti et al. (2004).
Proposition 3.2.
Suppose that $P_{0}$ is absolutely continuous with respect to the Lebesgue measure. Then there exists a random probability distribution $P_{\infty,x}$ such that $\lim_{N\to\infty}P_{N}(y)=P_{\infty,x}(y)=F_{\infty,x}(y)$ almost surely.
The proposition implies that the (random) limit $P_{\infty,x}$ is well defined and that the iterative sampling scheme is a valid way to draw from its distribution. We can be more precise about how fast this limit is approached.
Proposition 3.3.
For any $y\in\mathbb{R}$, $N\geq 0$, the following holds with probability at least $1-\delta$:
$|P_{\infty,x}(y)-P_{N}(y)|\lesssim\sqrt{\frac{\log(2/\delta)}{2\beta-1}}(n+N)^{-\beta+1/2}.$
The proof, given in Appendix A, is based on a time-uniform version of the Azuma-Hoeffding concentration inequality (Howard et al., 2020), the martingale property of the copula updates, and the learning rate schedule. Importantly, time-uniform martingale concentration allows us to directly bound the deviation of interest $|P_{\infty,x}-P_{N}|$ as opposed to $|P_{M}-P_{N}|$ with $M\geq N$ as in Fong et al. (2023).
Proposition 3.3 quantifies how well the distribution $P_{N}$ approximates $P_{\infty,x}$ after $N$ updates. Setting $N=0$ corresponds to the case where $P_{N}$ equals the initial PPD given only the observed data. In our setting, this is the output from the PFN. The random limit $P_{\infty,x}$ fluctuates around this point estimate, and the spread of these fluctuations characterizes the spread of the martingale posterior for $\theta(P(\cdot|x))$.
3.5 The Role of Initialization
To further understand the role of the initialization in Algorithm 1, the following theoretical statement can be made.
Corollary 3.1.
Let $P_{0,1}$ and $P_{0,2}$ be two PPDs to initialize our algorithm with, and $P_{\infty,x,1}$ and $P_{\infty,x,2}$ be the corresponding martingale limits after infinitely many iterations and $\Delta(y):=P_{\infty,x,1}(y)-P_{\infty,x,2}(y)$. Then for any $x,y$, it holds with probability at least $1-2\delta$ that
$|\Delta(y)|\geq|P_{0,1}(y)-P_{0,2}(y)|-O(n^{-\beta+1/2}).$
This is a direct result of Proposition 3.3 and shows that, up to the contraction rate of the posteriors, the difference in martingale posteriors is as large as the difference in the initializations. Hence, the choice of initialization is the dominating factor for the difference in the martingale posteriors.
3.6 The Role of $\beta$
The case $\beta=1$ corresponds to the choice $\alpha_{i}\sim i^{-1}$ proposed by Fong et al. (2023) in the unconditional context with contraction rate of $n^{-1/2}$. Contraction rates for nonparametric regression problems are typically much slower than $n^{-1/2}$ (Ghosal and Van der Vaart, 2017, Chapter 9), however. To accommodate such slow contraction rates, it is therefore crucial to allow for learning rates with $\beta<1$. The choice $\beta=d/(d+4)$ of our type 1 schedule corresponds to the optimal contraction rate we can expect in a nonparametric regression problem with $d$ features and twice continuously differentiable regression curve. However, Nagler (2023) demonstrated that current PFN architectures yield inconsistent estimators, as their bias does not vanish as $n\to\infty$. This suggests that to get valid inferences, the posterior should never contract fully. This happens in the regime $\beta\in(0,1/2]$, where the bound in Proposition 3.3 becomes vacuous, but the algorithm still converges to a random limit $P_{\infty,x}$ because of Proposition 3.2. This motivates the choice $\beta=4/(d+4)$, which has $\beta\leq 1/2$ for the regime $d\geq 4$, where the bias becomes notable already for moderate $n$.
3.7 Computation
In practice, we can only sample finite sequences and replace the MP by its finite approximation. The procedure is summarized in Algorithm 1. The overall run-time complexity of the algorithm is $O(BN)$, where $B$ is the number of replications, and $N$ is the length of one ‘chain’. Based on the considerations in the previous subsection, we recommend generating around $N=200$ forward samples in each replication. The computations are independent across the outer loop iterations ($b=1,\dots,B$) and straightforward to parallelize; the inner loop ($k=1,\dots,N$) must be run sequentially. Except for the initial computation of $\widehat{\operatorname{PPD}}(y\mid x)$, the runtime is independent of the dataset size $n$.
Algorithm 1 Computation of Martingale Posterior
1:Input: Estimated $\widehat{\operatorname{PPD}}(y|x)$ obtained from the PFN.
2:for $b=1,\dots,B$ do
3:  Initialize $P_{0}^{(b)}\leftarrow\widehat{\operatorname{PPD}}(y|x)$.
4:  for $k=1,\dots,N$ do
5:   Sample $y_{n+k}^{(b)}\sim P_{k-1}^{(b)}$.
6:   Update $(P_{k-1}^{(b)},y_{n+k}^{(b)})\to P_{k}^{(b)}$ as in (3).
7:  end for
8:  Compute $\widehat{P}_{N}^{(b)}(y)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Bigl\{y_{n+i}^{(b)}\leq y\Bigr\}$.
9:  Set $\theta^{(b)}(x)\leftarrow\theta\bigl(\widehat{P}_{N}^{(b)}\bigr)$.
10:end for
[…]
4 Numerical Experiments
To demonstrate the efficacy of our approach, we conduct various experiments. To estimate $\rho$ for DPMM-based methods, we always randomly draw $1000$ data points simulated by the PFN. If not specified explicitly, we use the default learning rate approach. We will refer to our method with this learning rate as Ours (default), while the adaptations proposed in Section 3.3 based on theoretical considerations and following the discussion about the biasedness of TabPFN are called Ours (type 1) and Ours (type 2). The original approach by Fong et al. (2023) is referred to as Fong et al. The approach that uses purely TabPFN for the forward process is called Pure TabPFN. Code is made available at https://anonymous.4open.science/r/UQ_for_PFNs.
4.1 TabPFN Martingale Property Check
In order to investigate whether TabPFN possesses the martingale property, we fit TabPFN using $n=25$ data points of a Gamma distribution with scale and shape parameter 2. Following the martingale routine in (1), we then predict $y_{n+k}$ for $k\in[1,1000]$ and always add the previous iteration’s sample to the dataset to refit TabPFN. In each step, we compute TabPFN’s PPD.
Results
As demonstrated in Figure 2, the resulting CDF of the iteratively updated PPD changes notably over time by first providing a wider PPD (all points of the QQ-plot lie below the bisecting line) and then becoming notably positively skewed for later iterations.
4.2 Convergence Speed
Figure 3: Effect of NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
We start by investigating the convergence speed of the MP sampling algorithm. For this, we subsample the concrete dataset (Yeh, 1998) to generate datasets of different sizes. After standardizing the $n\in\{25,100,250,1000\}$ data points, we visually inspect how the conditional posterior mean distribution of $y_{n+k}$ conditional on fixed features $x$ for $B=30$ instances evolves when running our MP approach for $k\in(0,1000]$ steps. Figure 3 summarizes the results for the default type learning rate and confirms that the different trajectories stabilize within the given budget of $N=1000$ sampling steps independent of $n$. This is confirmed in all of our practical applications, where convergence is observed no later than after $N=500$ steps. Results for different learning rate schedules are given in Section B.3.2.
4.3 Simulation Study
Our approach enables posterior estimation given new features $x$, complementing PFN’s aleatoric uncertainty with epistemic uncertainty. We demonstrate this using a simulation study based on Bayesian additive models. We generate the true underlying feature effects $f_{j}(x_{j})=B(x_{j})^{\top}\theta_{j}$ for feature $x_{j}$ by drawing parameters $\theta_{j}\sim\mathcal{N}(0,I_{20})$ and using those parameters as spline coefficients for a B-spline basis $B_{j}(x_{j})\in\mathbb{R}^{20}$ of degree 3 with 20 basis functions. The data is then generated by the law $Y|x,\theta\sim\mathcal{N}(\sum_{j=1}^{J}f_{j}(x_{j}),\sigma^{2})$ with mean determined by the additive model predictor and fixed variance $\sigma^{2}$. We use 30 features from which $\varphi\in\{1,3,10\}$ have a non-zero signal while the rest act as noise features with no contribution to $Y|x$. At the same time, we again only work with small to moderate amounts of training data $n\in\{30,100,300\}$ and fix $\sigma^{2}=0.5$. Due to the Gaussian likelihood and prior and the linearity of spline coefficients, the resulting posterior is analytically available and can be used to construct credible intervals for each of the functions $f_{j}$ (referred to as Optimal). This allows evaluating the quality of epistemic uncertainty. We use $N=250$ for all methods.
Table 1: Average conditional coverage across all $x$ locations for different methods (columns) for the simulation study using different $(n,\varphi)$ combinations (rows).
$n$ $\varphi$ Optimal Fong et al. Ours Pure TabPFN
30 1 0.91 (1.99) 0.93 (2.57) 0.90 (2.12) 0.83 (2.60)
30 3 0.74 (2.84) 0.65 (3.09) 0.88 (3.52) 0.87 (3.56)
30 10 0.82 (5.57) 0.99 (6.64) 0.94 (5.83) 0.74 (5.91)
100 1 0.92 (1.85) 0.85 (2.65) 0.89 (1.83) 0.82 (1.92)
100 3 0.93 (2.02) 0.96 (3.54) 0.92 (2.77) 0.87 (2.94)
100 10 0.99 (3.17) 0.05 (1.05) 0.99 (6.12) 0.94 (5.83)
300 1 0.91 (1.76) 0.58 (1.54) 0.90 (1.78) 0.84 (1.75)
300 3 0.91 (1.83) 0.92 (3.09) 0.90 (2.16) 0.67 (2.13)
300 10 0.95 (2.06) 0.93 (5.04) 0.97 (5.10) 0.94 (5.06)
Results
The results are depicted in Figure 1 and Table 1 while detailed results breaking down performance into different learning rates and feature locations can be found in Section B.2. While our approach is able to fit the data-generating process well and its uncertainty aligns with the ground truth posterior, initializing Algorithm 1 with the approach proposed Fong et al. (2023) yields a worse fit and coverage for the depicted function. This is expected as, already as a point estimator, Fong’s predictive is far worse than the state-of-the art performance that TabPFN delivers. The pure TabPFN forward process, which also initializes the PPD as in our approach, also shows worse performance and clearly demonstrates that the missing martingale property lets the PPD drift away from the initial estimate when using a pure TabPFN update instead of our proposal in Algorithm 1. Further note that, while all simulation settings, including our method with all three learning rates, can be run within 10 minutes on a conventional CPU laptop, the Pure TabPFN forward process requires more than 15 hours on a GeForce RTX 4070 Ti GPU.
4.4 Benchmark
Figure 4: Mean values and standard deviations (error bars) of miscoverage (observed minus nominal value; x-axis) and interval length (y-axis) of the 90% credible intervals produced by the original method (blue), our proposed adaptions (orange, purple), and the bootstrap, across the different benchmark datasets. The red vertical line indicates perfect coverage.
As an empirical evaluation, we compare our method using the previously presented learning rate approaches on real-world data from the UCI repository (Dua and Graff, 2017). We further compare against a bootstrap approach, which represents the natural comparison method, as it can also determine uncertainty for estimated parameters from PFN without additional parametric assumptions. In this benchmark, we run a conditional quantile regression ($\alpha\in\{0.5,0.9\}$) on the datasets airfoil ($n=1503$, $p=5$), boston ($n=252$, $p=15$), concrete ($n=1030$, $p=8$, Yeh, 1998), diabetes ($n=442$, $p=10$, Efron et al., 2004), energy ($n=768$, $p=8$, Tsanas and Xifara, 2012), and fish ($n=908$, $p=6$). These datasets provide an ideal benchmark for PFNs, which are still limited in their scalability to larger datasets. We use $n=100$ data points for training and aim to measure the uncertainty of PFN for the predicted quantile on the remaining data points. In addition to our approaches using the martingale posterior with $B=100$, the non-parametric bootstrap is computed. The bootstrap reruns the PFN method for $R=20$ times by drawing samples from the training dataset with replacement, makes quantile predictions, and computes the empirical 90% confidence interval of these predictions. We use both methods to construct a 90% credible/confidence interval (CI) and check the coverage of the predicted quantile given test dataset features. We also compute the length of these intervals to see how conservative our method is in comparison to a bootstrap approach, as well as the time for computation. We repeat the process five times with different splits.
Results
Figure 4 summarizes the results of our benchmark for $\alpha=0.5$, i.e., conditional median regression. The results are similar to those with smaller dataset size and for the quantile regression case (provided in Tables 8 and 9 in Appendix B), with our approach taking a fraction of the bootstrap approach’s time while yielding notably smaller CIs with better coverage. When comparing the different learning rate schemes, we see that in most cases, it is possible to either decrease the interval width (airfoil, boston, concrete, energy, fish) or improve the coverage (boston, concrete, energy, fish), or both, over Fong et al. (2023).
In addition, Section B.3.4 provides a comparison that fixes the time budget of both the MP and the bootstrap approach. As the bootstrap takes notably longer than our method (requiring $R$ forward passes of the PFN while our routine only uses PFNs once for $P_{0}$ and the estimation of $\rho$), the undercoverage of the bootstrap becomes even more pronounced.
Figure 5: Coverage and interval length behavior for the larger UCI datasets (rows) for differently sized training datasets.
4.5 Ablation: Dataset Size
In the previous sections, the sizes of the datasets are relatively small. Using small datasets is done intentionally, as PFNs particularly shine in small-scale applications by leveraging their pretraining. Here, we further investigate how coverages change when increasing the dataset size using the two largest datasets (airfoil, concrete) from Section 4.4 and subsample the datasets to training data sizes of $n\in\{50,100,200,500,700\}$.
Failure due to Bias
This analysis is particularly interesting in light of our discussion in Section 3.6, suggesting that intervals will not yield nominal coverage, even for $n\to\infty$ due to the non-vanishing bias of TabPFN. Empirically, this is validated in Figure 5, showing that neither the learning rate from Fong et al. (2023) nor our suggested alternative (type 1) obtains nominal coverage and even drifts away from the nominal level for increasing $n$.
In contrast, when using the type 2 approach suggested in Section 3.3, we find that our method is working as expected with decreasing interval lengths and decreasing deviations from the nominal coverage (0.9) for an increase in $n$.
4.6 Ablation: Initialization and Rollout
Initialization
To empirically validate that the initialization via PFN is the dominant factor in the learned posterior, we conduct another ablation study in which we swap the TabPFN initialization used in Section 4.4 with the PPD learned by the conditional copula regression method proposed in Fong et al. (2023). The learning-rate schedule, the bandwidth, how it is tuned, and the predictive resampling scheme are the same as in our previous experiments.
Rollout
Furthermore, we also investigate pure TabPFN updates of the PPD as suggested in Ng et al. (2025) using forward sampling of training features via Bayesian bootstrap and $N=1000$ forward samples. This ablation allows us to check the extent to which TabPFN’s missing martingale property discussed in Section 3 and the bias underlying the motivation for our proposals in Section 3.6 affect the posterior quality.
Results
The first three columns in Table 2 show the interval widths together with the coverages for our method and the three proposed $\beta$-schedules. While widths are in a similar range as the ones in Figure 4 with TabPFN initialization—which is to be expected as we control for all other tuning parameters—coverage values are far from the nominal level (90%). This is because the initial PPD by Fong et al. (2023) is a much worse estimator of the true conditional distribution than TabPFN, and confirms Corollary 3.1 and the importance of our TabPFN initialization.
The last column in Table 2 demonstrates that a PPD update purely based on TabPFN does not achieve nominal coverages either, even though the intervals are 5 to 10 times wider than the ones produced by our method. This is caused by interval centers drifting away from the initial estimate as a consequence of the lack of a martingale property.
Table 2: Coverage values with interval lengths in brackets across different datasets (rows) and methods.
Default Type 1 Type 2 Pure TabPFN
airfoil 0.18 (0.29) 0.05 (0.07) 0.12 (0.19) 0.27 (0.68)
boston 0.25 (0.23) 0.08 (0.08) 0.19 (0.18) 0.27 (0.51)
concrete 0.16 (0.30) 0.05 (0.08) 0.11 (0.21) 0.34 (1.02)
diabetes 0.34 (0.29) 0.12 (0.10) 0.27 (0.23) 0.65 (0.79)
energy 0.27 (0.12) 0.09 (0.04) 0.20 (0.09) 0.53 (0.89)
fish 0.24 (0.22) 0.06 (0.06) 0.17 (0.15) 0.54 (0.70)
5 Discussion
This work proposes an efficient and principled Bayesian uncertainty quantification method for estimates derived from prior fitted networks. Our method thereby provides epistemic uncertainty complementing the aleatoric uncertainty of PFNs. Epistemic uncertainty methods such as deep ensembles (Lakshminarayanan et al., 2017) are a potential alternative, but computationally expensive, as they require training multiple models. In the context of pretrained models, this is usually not a feasible option. In contrast, our method only retrains the martingale posterior routine, not the model itself, which is generally much more efficient.
Limitations
Some limitations remain in the proposed methodology. In particular, the current implementation does not offer credible sets that are uniform in the feature $x$, because there is no way to forward sample entire prediction curves. However, since our method is fast, allowing us to evaluate hundreds of different feature values within minutes, this is less of a problem with our method and more problematic for approaches such as the one from Ng et al. (2025) proposing pure PFN rollouts that can take hours for individual feature conditions. Our proposed methodology may also be applied to non-tabular or non-iid settings, but this is currently not covered by our theory and experiments. In particular, while our theory and algorithm can be transferred directly to other in-context predictors that provide a PPD, such as large language models (LLMs), this is only possible for certain types of prediction tasks, such as those proposed in Falck et al. (2024); Requeima et al. (2024). In this case, a text-based prompt is used to predict future numeric values, thereby reducing the large token space to, e.g., $[0,1]$ (Falck et al., 2024) or a 10-dimensional digit space (Requeima et al., 2024). To substantiate this hypothesis, we provide a proof-of-concept in Appendix D.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A PROOF OF PROPOSITION 3.3
We first show that $P_{M}(y)$ is a martingale. It holds
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=(1-\alpha_{n+M})P_{M-1}(y)+\alpha_{n+M}\mathbb{E}_{y_{n+M}\sim P_{M-1}}[H_{\rho}(P_{M-1}(y),P_{M-1}(y_{n+M})].$
By the probability integral transform, it holds $P_{M}(y_{n+M+1})\sim\mathrm{Uniform}[0,1]$. Thus,
$\mathbb{E}_{y_{n+M}\sim P_{M-1}}[H_{\rho}(P_{M-1}(y),P_{M}(y_{n+M})]=\int H_{\rho}(P_{M-1}(y),u)du=P_{M-1}(y),$
by the properties of the Gaussian copula. Hence,
$\mathbb{E}[P_{M}(y)|y_{(n+1):(n+M-1)}]=P_{M-1}(y),$
which implies that $P_{M}(y)$ is a martingale adapted to the filtration $\sigma(y_{1:n})_{n\in\mathbb{N}}$.
Next, we apply the uniform-in-time version of the Hoeffding-Azuma inequality given in Corollary 1 (a) of Howard et al. (2020). It holds
$|P_{M}(y)-P_{M-1}(y)|\leq\alpha_{n+M}\leq C(n+M+1)^{-\beta},\quad\text{for all }y\in\mathbb{R},$
and, for all $M\geq N$,
$\sum_{i=N}^{M}\alpha_{n+i}^{2}\leq\sum_{i=N}^{\infty}\alpha_{n+i}^{2}\leq C^{2}\int_{N}^{\infty}(n+t)^{-2\beta}dt\leq C_{\beta}(n+N)^{1-2\beta},$
where $C_{\beta}=C^{2}/(2\beta-1)$.
Now Corollary 1 (a) of Howard et al. (2020) with $m=C_{\beta}(n+N)^{1-2\beta}$ yields
$\Pr(\exists M\geq N\colon|P_{M}(y)-P_{N}(y)|\geq\epsilon)\leq 2\exp\left(-\frac{\epsilon^{2}}{2m}\right)=2\exp\left(-\frac{\epsilon^{2}(n+N)^{2\beta-1}}{2C_{\beta}}\right).$
Setting ${\epsilon}=\sqrt{C_{\beta}\log(2/\delta)}(n+N)^{1/2-\beta}$ gives,
$\Pr(\exists M\geq N\colon|P_{M}(y)-P_{N}(y)|\geq\epsilon)\leq\delta.$
Now the claim follows since $P_{M}(y)\to P_{\infty,x}(y)$ almost surely (Proposition 3.2). ∎
Appendix B Further Numerical Experiments and Details
B.1 Estimation of $\rho$
We implemented the estimation of $\rho$ for the Gaussian copula update as follows: If the computational budget is limited, we follow the routine suggested by Fong et al. (2023) and maximize the prequential log score $\sum_{i=1}^{n}\log p_{i-1}(y_{i}|\bm{x}_{i})$ for $\rho$ (an efficient method is implemented in the MP module by the same authors). If the dataset size is limited or if one likes to separate tuning from training, one could potentially draw a new and larger set of density values from TabPFN and independently estimate using this “tuning set”. In practice, we found the training data to be sufficient for good estimation of $\rho$. We also reuse this estimator for all predictive steps in our Martingale Posterior computation and have not found any significant improvement by re-estimating the correlation or manually tuning via cross-validation.
B.2 Additional Simulation Results
The details for the conditional coverages for all simulation settings of the study described in Section 4.3 can be found in Tables 3, 4 and 5
Table 3: Detailed simulation results for $n=30$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.716 (2.113) 0.963 (2.588) 0.934 (2.219) 0.942 (2.254) 0.942 (2.254) 0.998 (3.152)
-1.94 0.947 (1.990) 0.947 (2.588) 0.943 (2.395) 0.949 (2.430) 0.949 (2.430) 0.945 (2.007)
-1.39 0.961 (2.148) 0.990 (2.588) 0.974 (2.307) 0.970 (2.307) 0.970 (2.307) 1.000 (5.846)
-0.83 0.960 (2.113) 0.956 (2.588) 0.931 (2.254) 0.945 (2.307) 0.945 (2.307) 0.985 (2.641)
-0.28 0.920 (1.831) 0.986 (2.571) 0.978 (2.377) 0.977 (2.395) 0.977 (2.395) 0.910 (1.902)
0.28 0.845 (1.937) 0.965 (2.571) 0.959 (2.342) 0.963 (2.360) 0.963 (2.360) 0.920 (2.148)
0.83 0.938 (1.831) 0.943 (2.571) 0.861 (2.360) 0.855 (2.360) 0.855 (2.360) 0.902 (2.024)
1.39 0.925 (1.867) 0.989 (2.571) 0.976 (2.272) 0.976 (2.289) 0.976 (2.289) 0.788 (2.258)
1.94 0.941 (1.955) 0.604 (2.553) 0.589 (2.324) 0.604 (2.324) 0.604 (2.324) 0.564 (2.208)
2.50 0.910 (2.095) 0.939 (2.553) 0.921 (2.272) 0.938 (2.324) 0.938 (2.324) 0.967 (2.623)
3 -2.50 0.285 (2.787) 0.906 (3.612) 0.906 (3.592) 0.906 (3.592) 0.906 (3.592) 0.998 (7.830)
-1.94 0.287 (2.787) 0.910 (3.612) 0.896 (3.651) 0.896 (3.690) 0.896 (3.690) 0.947 (3.442)
-1.39 0.987 (2.945) 0.999 (3.612) 0.998 (3.494) 0.999 (3.553) 0.999 (3.553) 1.000 (4.072)
-0.83 0.869 (2.905) 0.989 (3.455) 0.995 (3.533) 0.995 (3.553) 0.995 (3.553) 0.936 (2.280)
-0.28 0.893 (2.552) 0.441 (2.925) 0.900 (3.514) 0.900 (3.533) 0.900 (3.533) 0.697 (3.600)
0.28 0.774 (2.768) 0.001 (1.963) 0.748 (3.533) 0.748 (3.553) 0.748 (3.553) 0.573 (3.090)
0.83 0.731 (2.630) 0.087 (1.924) 0.989 (3.514) 0.987 (3.592) 0.987 (3.592) 0.973 (7.290)
1.39 0.993 (3.239) 0.782 (2.768) 0.995 (3.416) 0.995 (3.416) 0.995 (3.416) 0.960 (2.878)
1.94 0.893 (3.062) 0.856 (3.416) 0.942 (3.514) 0.942 (3.553) 0.942 (3.553) 0.794 (2.696)
2.50 0.651 (2.729) 0.500 (3.612) 0.439 (3.455) 0.439 (3.475) 0.439 (3.475) 0.771 (4.372)
10 -2.50 0.977 (5.679) 1.000 (6.659) 1.000 (6.017) 1.000 (6.085) 1.000 (6.085) 0.970 (6.390)
-1.94 0.970 (5.679) 1.000 (6.659) 0.998 (5.949) 0.998 (5.983) 0.998 (5.983) 0.948 (5.401)
-1.39 1.000 (5.713) 1.000 (6.659) 1.000 (6.051) 1.000 (6.118) 1.000 (6.118) 1.000 (6.645)
-0.83 1.000 (5.442) 1.000 (6.659) 1.000 (5.882) 1.000 (5.882) 1.000 (5.882) 0.982 (5.352)
-0.28 0.865 (5.307) 0.996 (6.659) 0.936 (6.118) 0.967 (6.220) 0.967 (6.220) 0.773 (5.792)
0.28 0.974 (5.476) 1.000 (6.659) 1.000 (6.186) 1.000 (6.186) 1.000 (6.186) 0.995 (6.247)
0.83 1.000 (5.713) 1.000 (6.625) 1.000 (6.118) 1.000 (6.220) 1.000 (6.220) 0.969 (6.250)
1.39 0.678 (5.442) 0.998 (6.625) 0.968 (5.611) 0.968 (5.611) 0.968 (5.611) 0.683 (5.014)
1.94 0.431 (5.645) 0.953 (6.558) 0.842 (5.848) 0.842 (5.780) 0.842 (5.780) 0.426 (5.385)
2.50 0.340 (5.611) 0.922 (6.592) 0.831 (6.186) 0.865 (6.220) 0.865 (6.220) 0.668 (5.842)
Table 4: Detailed simulation results for $n=100$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.959 (2.010) 0.998 (3.134) 0.945 (1.912) 0.945 (1.912) 0.945 (1.912) 0.964 (2.247)
-1.94 0.902 (1.853) 0.786 (3.075) 0.801 (1.675) 0.801 (1.675) 0.801 (1.675) 0.766 (1.261)
-1.39 0.922 (1.754) 0.990 (2.937) 0.890 (1.596) 0.890 (1.596) 0.890 (1.596) 0.927 (2.148)
-0.83 0.898 (1.794) 0.977 (2.621) 0.863 (1.557) 0.863 (1.557) 0.863 (1.557) 0.917 (1.872)
-0.28 0.888 (1.715) 0.475 (2.247) 0.892 (1.675) 0.892 (1.675) 0.892 (1.675) 0.770 (1.202)
0.28 0.913 (1.715) 0.923 (2.050) 0.902 (1.794) 0.902 (1.794) 0.902 (1.794) 0.922 (1.899)
0.83 0.924 (1.754) 0.379 (2.129) 0.935 (1.813) 0.935 (1.813) 0.935 (1.813) 0.941 (1.907)
1.39 0.951 (1.931) 0.983 (2.385) 0.986 (2.385) 0.986 (2.385) 0.986 (2.385) 0.924 (2.060)
1.94 0.900 (1.774) 0.986 (2.818) 0.856 (1.636) 0.856 (1.636) 0.856 (1.636) 0.907 (1.895)
2.50 0.947 (2.148) 0.997 (3.075) 0.932 (1.872) 0.932 (1.872) 0.932 (1.872) 0.991 (2.591)
3 -2.50 0.976 (2.243) 0.998 (3.602) 0.745 (2.786) 0.745 (2.786) 0.745 (2.786) 0.993 (5.536)
-1.94 0.902 (2.039) 0.950 (3.579) 0.978 (2.696) 0.978 (2.696) 0.978 (2.696) 0.998 (6.488)
-1.39 0.945 (1.948) 0.861 (3.579) 0.935 (2.673) 0.935 (2.673) 0.935 (2.673) 0.923 (2.512)
-0.83 0.905 (1.948) 0.998 (3.602) 0.985 (2.900) 0.985 (2.900) 0.985 (2.900) 0.976 (3.209)
-0.28 0.921 (1.880) 0.885 (3.443) 0.992 (3.013) 0.992 (3.013) 0.992 (3.013) 0.860 (2.647)
0.28 0.937 (1.835) 0.980 (3.239) 0.995 (2.786) 0.995 (2.786) 0.995 (2.786) 0.876 (2.466)
0.83 0.902 (1.903) 0.970 (3.579) 0.996 (2.990) 0.996 (2.990) 0.996 (2.990) 0.748 (2.433)
1.39 0.913 (1.926) 0.952 (3.602) 0.981 (2.786) 0.981 (2.786) 0.981 (2.786) 0.938 (3.049)
1.94 0.926 (2.152) 0.998 (3.602) 0.850 (2.718) 0.850 (2.718) 0.850 (2.718) 0.878 (2.461)
2.50 0.979 (2.288) 0.997 (3.602) 0.997 (2.877) 0.997 (2.877) 0.997 (2.877) 0.929 (2.588)
10 -2.50 0.980 (3.219) 0.004 (1.048) 0.999 (6.214) 0.999 (6.214) 0.999 (6.214) 0.996 (4.901)
-1.94 0.998 (3.294) 0.318 (1.048) 1.000 (6.289) 1.000 (6.289) 1.000 (6.289) 1.000 (7.105)
-1.39 0.992 (3.294) 0.006 (1.048) 1.000 (6.476) 1.000 (6.476) 1.000 (6.476) 0.999 (6.446)
-0.83 0.998 (3.107) 0.009 (1.048) 1.000 (6.214) 1.000 (6.214) 1.000 (6.214) 0.997 (6.578)
-0.28 0.997 (3.069) 0.004 (1.048) 1.000 (6.363) 1.000 (6.363) 1.000 (6.363) 0.992 (5.660)
0.28 0.996 (3.032) 0.001 (1.048) 0.999 (6.588) 0.999 (6.625) 0.999 (6.625) 0.950 (5.060)
0.83 0.991 (3.107) 0.005 (1.048) 1.000 (6.326) 1.000 (6.326) 1.000 (6.326) 0.943 (5.035)
1.39 0.997 (3.069) 0.094 (1.048) 1.000 (6.251) 1.000 (6.251) 1.000 (6.251) 0.969 (5.869)
1.94 0.999 (3.219) 0.038 (1.048) 1.000 (5.802) 1.000 (5.802) 1.000 (5.802) 0.802 (5.548)
2.50 0.924 (3.294) 0.000 (1.048) 0.995 (6.513) 0.995 (6.513) 0.995 (6.513) 0.701 (5.039)
Table 5: Detailed simulation results for $n=300$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.954 (2.020) 0.935 (1.827) 0.926 (1.770) 0.926 (1.770) 0.926 (1.770) 0.982 (2.501)
-1.94 0.922 (1.866) 0.892 (1.770) 0.884 (1.770) 0.884 (1.770) 0.884 (1.770) 0.869 (1.520)
-1.39 0.894 (1.674) 0.613 (1.674) 0.879 (1.577) 0.879 (1.577) 0.879 (1.577) 0.861 (1.789)
-0.83 0.898 (1.674) 0.443 (1.577) 0.915 (1.750) 0.915 (1.750) 0.915 (1.750) 0.907 (1.712)
-0.28 0.907 (1.674) 0.756 (1.539) 0.835 (1.404) 0.835 (1.404) 0.835 (1.404) 0.698 (1.366)
0.28 0.903 (1.674) 0.502 (1.481) 0.897 (1.635) 0.897 (1.635) 0.897 (1.635) 0.932 (1.904)
0.83 0.899 (1.674) 0.793 (1.423) 0.951 (2.231) 0.951 (2.231) 0.951 (2.231) 0.975 (2.526)
1.39 0.898 (1.693) 0.780 (1.385) 0.954 (2.154) 0.954 (2.154) 0.954 (2.154) 0.961 (2.087)
1.94 0.922 (1.750) 0.101 (1.347) 0.932 (1.847) 0.932 (1.847) 0.932 (1.847) 0.957 (1.928)
2.50 0.920 (1.943) 0.001 (1.327) 0.909 (2.001) 0.909 (2.001) 0.909 (2.001) 0.783 (1.676)
3 -2.50 0.924 (2.113) 0.860 (3.090) 0.960 (2.681) 0.960 (2.681) 0.960 (2.681) 0.977 (3.849)
-1.94 0.927 (1.818) 0.945 (3.090) 0.776 (2.045) 0.776 (2.045) 0.776 (2.045) 0.950 (3.560)
-1.39 0.916 (1.727) 0.864 (3.090) 0.905 (1.863) 0.905 (1.863) 0.905 (1.863) 0.886 (2.920)
-0.83 0.915 (1.727) 0.950 (3.090) 0.948 (2.045) 0.948 (2.045) 0.948 (2.045) 0.934 (2.777)
-0.28 0.911 (1.727) 0.967 (3.090) 0.967 (2.181) 0.967 (2.181) 0.967 (2.181) 0.890 (2.365)
0.28 0.904 (1.704) 0.955 (3.090) 0.962 (2.158) 0.962 (2.158) 0.962 (2.158) 0.875 (2.817)
0.83 0.887 (1.749) 0.998 (3.090) 0.922 (2.158) 0.922 (2.158) 0.922 (2.158) 0.837 (2.542)
1.39 0.907 (1.727) 0.996 (3.090) 0.985 (2.499) 0.985 (2.499) 0.985 (2.499) 0.926 (2.900)
1.94 0.925 (1.795) 0.982 (3.090) 0.959 (2.158) 0.959 (2.158) 0.959 (2.158) 0.969 (3.095)
2.50 0.891 (2.181) 0.708 (3.090) 0.694 (2.545) 0.694 (2.545) 0.694 (2.545) 0.671 (2.032)
10 -2.50 0.977 (2.298) 0.988 (6.321) 1.000 (5.377) 1.000 (5.377) 1.000 (5.377) 0.999 (6.187)
-1.94 0.934 (2.052) 0.401 (5.869) 0.880 (5.377) 0.880 (5.377) 0.880 (5.377) 0.880 (5.085)
-1.39 0.940 (2.011) 1.000 (5.418) 1.000 (5.089) 1.000 (5.089) 1.000 (5.089) 1.000 (5.382)
-0.83 0.951 (2.011) 1.000 (4.966) 1.000 (5.171) 1.000 (5.171) 1.000 (5.171) 1.000 (5.186)
-0.28 0.931 (1.929) 0.997 (4.474) 1.000 (5.582) 1.000 (5.582) 1.000 (5.582) 0.979 (5.313)
0.28 0.944 (1.970) 1.000 (4.227) 1.000 (5.089) 1.000 (5.089) 1.000 (5.089) 0.957 (5.104)
0.83 0.945 (1.929) 0.943 (4.227) 1.000 (5.541) 1.000 (5.541) 1.000 (5.541) 0.973 (5.684)
1.39 0.947 (1.970) 1.000 (4.515) 1.000 (5.418) 1.000 (5.418) 1.000 (5.418) 0.983 (5.215)
1.94 0.955 (2.052) 0.997 (4.925) 1.000 (5.254) 1.000 (5.254) 1.000 (5.254) 0.995 (5.615)
2.50 0.982 (2.422) 0.997 (5.500) 1.000 (5.336) 1.000 (5.336) 1.000 (5.336) 0.998 (6.050)
B.3 Additional Experiments
B.3.1 Conditional Quantile Coverages
While the previous example evaluates the obtained posterior as a whole, we now investigate its coverage properties for conditional statistics. For this, we simulate data using a heteroscedastic Gaussian funnel $\mathcal{N}(\sin(3x),x^{2})$ for $x\in[0,1]$ for $n=100$ data points. We then compute the posterior 90%-credible intervals (CIs) for the quantiles $P^{-1}(\alpha|x,\mathcal{D}_{n})$, various $\alpha$-levels, and three values of interest $x\in\{0.2,0.5,0.8\}$. For our routine, we use $B=400$. We repeat this process 20 times with a new random dataset and check how often the true quantile is included in the computed CI in each repetition.
Table 6: Coverages ($\pm$ two std. errors) of different quantiles (columns) at different $x$ values (rows), highlighted where ranges include the nominal coverage.
$x\backslash\alpha$ 0.05 0.10 0.25 0.5 0.75 0.90 0.95
$x=0.2$ 0.65 $\pm$ 0.21 0.60 $\pm$ 0.22 0.65 $\pm$ 0.21 0.75 $\pm$ 0.19 0.80 $\pm$ 0.18 0.85 $\pm$ 0.16 0.85 $\pm$ 0.16
$x=0.5$ 0.70 $\pm$ 0.21 0.70 $\pm$ 0.21 0.75 $\pm$ 0.19 0.85 $\pm$ 0.16 0.90 $\pm$ 0.13 0.75 $\pm$ 0.19 0.70 $\pm$ 0.21
$x=0.8$ 0.95 $\pm$ 0.10 0.85 $\pm$ 0.16 0.90 $\pm$ 0.13 1.00 $\pm$ 0.00 0.85 $\pm$ 0.16 0.70 $\pm$ 0.21 0.65 $\pm$ 0.21
Results
The data is visualized in Figure 9 in Appendix B. The resulting coverages are shown in Table 6. Our method provides (close to) nominal coverage for many combinations of $x$ and $\alpha$ but is less accurate for $x=0.2$ and extreme quantiles in general. This can be explained by the fact that the data at $x=0.2$ has almost no variation, while little data is available for extreme values, making these cases very challenging given only 100 data points.
Bootstrap Comparison
To show that this is a difficult task, we further provide results for the bootstrap in Table 7. The results show that the bootstrap (with $B=400$) also struggles to obtain nominal coverages, and for all larger x- and more extreme quantile values, with significantly lower coverages than 90%.
Table 7: Results for different values of $x$ and $\alpha$.
$x\backslash\alpha$ 0.05 0.10 0.25 0.50 0.75 0.90 0.95
0.2 0.65 $\pm$ 0.30 0.70 $\pm$ 0.29 0.70 $\pm$ 0.29 0.80 $\pm$ 0.25 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.65 $\pm$ 0.30
0.5 0.70 $\pm$ 0.29 0.85 $\pm$ 0.26 0.65 $\pm$ 0.30 0.75 $\pm$ 0.28 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.65 $\pm$ 0.30
0.8 0.40 $\pm$ 0.31 0.55 $\pm$ 0.32 0.75 $\pm$ 0.28 0.80 $\pm$ 0.25 0.70 $\pm$ 0.29 0.50 $\pm$ 0.32 0.60 $\pm$ 0.31
B.3.2 Convergence Speed: Additional Results
Figures 6 and 7 depict the convergence speeds for the analysis in Section 4.2 for the two other learning rate schedules, type 1 and type 2, respectively. In this example, all results look qualitatively very similar.
Figure 6: Effect of number of generated samples NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
Figure 7: Effect of number of generated samples NN on the convergence of $P_{N}$, depicted by $B=30$ different PPD mean estimates.
B.3.3 Benchmarks: Additional Results
Tables 8 and 9 summarize the results of our benchmark for $\alpha=0.5$ and $\alpha=0.9$, respectively for $n=25$. While both methods often have very close to nominal coverage or provide larger than nominal coverage, the bootstrap method fails in one case to provide a CI with close to nominal coverage. Despite having larger than nominal coverage, our MP approach consistently yields shorter or equally large CIs while often being twice as fast as the bootstrap.
Table 8: Comparison of the Martingale Posterior and Bootstrap method across datasets for conditional median regression ($\alpha=0.5$). Reported are mean $\pm$ standard deviation for coverage, interval length, and runtime (seconds).
Martingale Posteriors ($B=100)$ Bootstrap ($R=20)$
Coverage Interval Length Time Coverage Interval Length Time
airfoil 0.96 $\pm$ 0.01 0.23 $\pm$ 0.04 153.39 $\pm$ 29.31 0.91 $\pm$ 0.01 0.44 $\pm$ 0.03 328.97 $\pm$ 18.12
boston 0.93 $\pm$ 0.01 0.25 $\pm$ 0.02 267.72 $\pm$ 27.62 0.98 $\pm$ 0.02 0.51 $\pm$ 0.03 209.23 $\pm$ 13.93
concrete 0.93 $\pm$ 0.01 0.32 $\pm$ 0.05 125.38 $\pm$ 13.19 0.74 $\pm$ 0.03 0.45 $\pm$ 0.02 324.86 $\pm$ 17.52
diabetes 0.96 $\pm$ 0.03 0.70 $\pm$ 0.12 260.96 $\pm$ 14.15 0.94 $\pm$ 0.02 1.00 $\pm$ 0.06 159.72 $\pm$ 28.37
energy 0.97 $\pm$ 0.00 0.09 $\pm$ 0.02 294.70 $\pm$ 24.77 0.83 $\pm$ 0.09 0.09 $\pm$ 0.01 232.36 $\pm$ 14.36
fish 0.95 $\pm$ 0.02 0.38 $\pm$ 0.05 113.53 $\pm$ 17.79 0.94 $\pm$ 0.02 0.69 $\pm$ 0.05 212.40 $\pm$ 27.32
Table 9: Comparison of the Martingale Posterior and Bootstrap method across datasets for conditional quantile regression ($\alpha=0.9$). Reported are mean $\pm$ standard deviation for coverage, interval length, and runtime (seconds).
Martingale Posteriors ($B=100)$ Bootstrap ($R=20)$
Coverage Interval Length Time Coverage Interval Length Time
airfoil 0.96 $\pm$ 0.01 0.25 $\pm$ 0.02 164.01 $\pm$ 11.76 0.89 $\pm$ 0.01 0.52 $\pm$ 0.03 340.42 $\pm$ 19.23
boston 0.90 $\pm$ 0.01 0.29 $\pm$ 0.02 251.48 $\pm$ 22.59 0.96 $\pm$ 0.01 0.74 $\pm$ 0.11 187.17 $\pm$ 24.83
concrete 0.88 $\pm$ 0.02 0.32 $\pm$ 0.03 145.49 $\pm$ 19.84 0.65 $\pm$ 0.05 0.56 $\pm$ 0.03 382.44 $\pm$ 80.95
diabetes 0.84 $\pm$ 0.04 0.61 $\pm$ 0.13 296.40 $\pm$ 41.68 0.92 $\pm$ 0.02 1.00 $\pm$ 0.12 238.02 $\pm$ 67.33
energy 0.96 $\pm$ 0.01 0.12 $\pm$ 0.04 101.21 $\pm$ 16.50 0.74 $\pm$ 0.06 0.14 $\pm$ 0.03 226.96 $\pm$ 17.62
fish 0.87 $\pm$ 0.03 0.45 $\pm$ 0.08 100.80 $\pm$ 28.50 0.94 $\pm$ 0.01 0.87 $\pm$ 0.13 211.07 $\pm$ 25.82
B.3.4 Benchmarks: Comparison Using Fixed Time Budget
To make a fair comparison between our method and the bootstrap, we run additional benchmarks with an equal time budget for both methods. The results are given in Table 10. These results further strengthen the case for our method by showing that for equal time budget, the Bootstrap produces much wider intervals that nevertheless fail to reach the target coverage in several cases.
Table 10: Coverage and interval lengths for MP and Bootstrap methods.
Dataset Coverage (MP) Interval Length (MP) Coverage (Bootstrap) Interval Length (Bootstrap)
airfoil 0.96 $\pm$ 0.01 0.23 $\pm$ 0.05 0.87 $\pm$ 0.03 0.40 $\pm$ 0.04
boston 0.90 $\pm$ 0.01 0.25 $\pm$ 0.02 0.92 $\pm$ 0.06 0.42 $\pm$ 0.03
concrete 0.91 $\pm$ 0.01 0.33 $\pm$ 0.04 0.64 $\pm$ 0.03 0.38 $\pm$ 0.03
diabetes 0.90 $\pm$ 0.03 0.65 $\pm$ 0.16 0.82 $\pm$ 0.07 0.82 $\pm$ 0.13
energy 0.97 $\pm$ 0.01 0.09 $\pm$ 0.01 0.73 $\pm$ 0.10 0.08 $\pm$ 0.01
fish 0.92 $\pm$ 0.02 0.38 $\pm$ 0.05 0.91 $\pm$ 0.05 0.64 $\pm$ 0.06
B.4 Conditional Regression Posterior
Refer to caption Refer to caption Refer to caption
Figure 8: Visualization of the diffusion process data (left) together with the PFN predicted density (center) and the estimated epistemic uncertainty (right) via our martingale posterior approach.
Our approach enables posterior estimation given new features $x$, complementing PFN’s aleatoric uncertainty with epistemic uncertainty. We demonstrate this using a diffusion process ($n=200$). This data is challenging as it evolves from an unimodal to a trimodal heteroscedastic Gaussian distribution with linear trends and sinusoidal changes depending on $x$. We run our routine and compute both the estimated density by TabPFN and the variance of this density as a measure of epistemic uncertainty provided by our method.
Results
The results are depicted in Figure 8. While TabPFN is able to fit the data-generating process well and its uncertainty (normalized density values) aligns with the apparent data for larger $x$ values, higher density values also coincide with higher uncertainty (as measured by the normalized density variance). This becomes particularly visible in the lower branch of the diffusion process, where less training data is available. The provided density by TabPFN is comparatively small in this area, while the corresponding epistemic uncertainty is relatively large.
B.5 Details on Data Generation
B.5.1 Data Generation for Diffusion Process
For the diffusion process data in Section B.4, we use the following data-generating process:
  1. 1.
    Input generation: The input variable $x$ is sampled uniformly from $[2.5,12.5]$.
  2. 2.
    Piecewise functional behavior: The output $y$ is determined by three different functions applied to $x$, separated by a data-dependent midpoint $m$, approximately at $\text{median}(x)-2$. The three functional forms are:
    • $f_{1}(x)$: A linear and sinusoidal function switching at $m$.
    • $f_{2}(x)$: A mirrored version of $f_{1}(x)$.
    • $f_{3}(x)$: A piecewise function with a constant component and a high-frequency sine term.
    Each sample is randomly assigned one of these three functions.
  3. 3.
    Heteroscedastic noise: Additive noise is introduced with a scale that increases quadratically with $x$, making uncertainty larger for larger $x$.
  4. 4.
    Output transformation: The input is centered to be within [2.5,2.5][-2.5,2.5], and the output values are normalized based on their min-max range for potential use in quantile-based learning.
B.5.2 Data Generation for Quantile Regression
For the data in Section B.3.1, we sample values $x\sim U(0,1)$ and generate the corresponding outcome via $y\sim\mathcal{N}(\sin(3x),x^{2})$. A visualization of the dataset is given in Figure 9.
Figure 9: Depiction of the data used in Section B.3.1.
B.6 Computational Environment
All computations except for the pure TabPFN rollout were performed on a user PC with Intel(R) Core(TM) i7-8665U CPU @ 1.90GHz, 8 cores, 16 GB RAM, using Python 3.12.7. Run times of each experiment do not exceed 24 hours. For the TabPFN rollout, computations were done on a GeForce RTX 4070 Ti GPU, with every single experiment exceeding 1 hour of runtime. All experiments using TabPFN are run with tabpfn==2.0.5.
GPT 4o and GPT o3-mini-high are called via OpenAI’s UI. DeepSeek R1 is called from its UI at www.deepseek.com. All other methods are implemented from scratch by the authors.
Appendix C Further Discussion
Another open problem is joint posterior inference for a collection of parameters $\theta=\{\theta(x)\colon x\in\mathcal{X}\}$. Although our proposed inference procedure can be repeated for many values of $x$, the resulting posteriors are disconnected. For obtaining a full joint posterior $\Pi(\theta|\mathcal{D}_{n})$, the distribution of the features $x_{1:n}$ can no longer be ignored. Fong et al. (2023) proposed a general joint update of the PPDs for all values of $x$ simultaneously. However, this general update is intractable, and the heuristic simplifications proposed by Fong et al. (2023) are neither particularly simple nor theoretically justified. There are several potential ways forward. A simple practical solution would be to specify a joint distribution that combines the individual posteriors $P_{\infty,x_{1}},\dots,P_{\infty,x_{K}}$ in a plausible way; for example, using a multivariate Gaussian copula with covariance kernel depending on the distance between values $x_{i}\neq x_{j}$. We expect such a heuristic correction to work reasonably well in many applications. A more sophisticated alternative was recently proposed by Huk et al. (2024) and involves nonparametric estimation of the implicit dependence between PPDs by a nonparametric vine copula.
Appendix D LLM Proof-of-Concept
To substantiate our previous hypothesis, we provide a proof-of-concept by prompting different LLMs to construct a CDF based on $n=25$ using the setup of Falck et al. (2024). The prompt is given in the following:
Given the following data which are i.i.d. sampled from a Gaussian,
-1.609 -0.384 -0.592  0.073 -1.998 -1.093 -1.033 -3.159 -1.714 -2.754  0.832
1.294  0.411 -1.168 -1.246 -1.279 -1.488  0.376 -1.554 -0.318  1.267  0.836
-1.456 -1.921 -2.545
generate the CDF of the posterior predictive for a new value,
evaluated at the following data points:
-5 -4.75 -4.5 -4.25 -4 -3.75 -3.5 -3.25 -3 -2.75 -2.5 -2.25 -2 -1.75 -1.5
-1.25 -1 -0.75 -0.5 -0.25 0 0.25 0.5 0.75 1 1.25 1.5 1.75 2 2.25 2.5 2.75 3
While Falck et al. (2024) investigated whether different LLMs possess the martingale property, we are interested in whether our approach can provide uncertainty for these models. To do so, we simulate i.i.d. data from a Gaussian distribution and provide it to the LLM (GPT-4o, GPT-o3-mini-high, DeepSeek R1) including a context that specifies how the data was generated. We then ask the LLM to generate CDF values $F(y_{i})$ for this distribution for a prespecified grid of $y_{i}$ values. Figure 10 shows the initial output CDF of the LLM as a red dashed line. The underlying true CDF is marked with a dashed-dotted green line, indicating that GPT-4o is unable to predict the CDF, while the others do so correctly. We then run $N=1000$ replications of forward loops of our Algorithm 1 to sample values $y_{(n+1):(n+N)}^{(b)}$ and plot the resulting CDF lines for $y_{1:(n+N)}^{(b)},b=1,\ldots,B$ in blue. These lines indicate the uncertainty of the CDF generated by the LLM. The average CDF resulting from these blue curves is plotted in yellow, indicating that all models have the martingale property (although GPT-4o is not correct in its prediction).
While the blue curves also nicely depict the epistemic uncertainty of the LLM for the predicted CDFs, it becomes clear that the provided method—as expected—cannot counteract biases in the LLM (as is the case for GPT 4o). For more advanced LLMs, however, our approach gives uncertainty quantification akin to what is investigated Requeima et al. (2024), but potentially much cheaper.
Figure 10: Predicted CDFs using different LLMs (facets) with MP uncertainty measure visualized in blue using $B=1000$ replications and $N=1000$ samples.
Posteriors for LLM-Generated Text
PFNs are in-context learners based on the transformer architecture, just as modern LLMs. LLMs are trained to approximate the PPD $p(y|x,\mathcal{D})$ of the next token $y$ given a context $x$ of previous tokens and the corpus $\mathcal{D}$ they were trained on. Given this similarity, the martingale posterior approach could potentially also be adapted to quantify uncertainty in LLM-generated text. There are, however, two major obstacles. First, simple PPD updates, as in (3), are impossible due to the complex relation between the next and the previous tokens. The complexity of such updates can likely only be matched by an equally strong LLM. Second, Proposition 3.1 and Proposition 3.2 fail due to the lack of exchangeability in the generated output tokens. Contraction results like Proposition 3.3, however, rely on only the martingale property and would remain valid, although the interpretation of the distributions $P_{M}$ for $M\to\infty$ is less clear. Investigating how and to what extent martingale posterior ideas transfer to language models is an exciting path for future research.